# RobustVLA: Robustness-Aware Reinforcement Post-Training for Vision-Language-Action Models

### v1

@@ 1 Introduction
Vision-Language-Action (VLA) models have emerged as a powerful paradigm for general-purpose robotic manipulation, with representative examples including RT-1 (Brohan et al., 2022), RT-2 (Zitkovich et al., 2023), OpenVLA (Kim et al., 2024), and $\pi_{0}$ (Black et al., 2024). These VLA models demonstrate the feasibility of learning universal robotic policies from large-scale multimodal datasets and have achieved remarkable performance across a variety of manipulation tasks. Despite these advances, recent studies (Xing et al., 2025; Shi et al., 2025) reveal that the generalization ability of pre-trained VLA models remains limited when deployed in Out-of-Distribution (OOD) scenarios. VLA models often struggle with unseen objects, novel environments, and different robot embodiments. Even minor changes in object placement or environmental conditions, such as lighting variations, can significantly degrade performance. One key reason is that VLA models may rely on spurious correlations between irrelevant observation features and actions, rather than capturing the true causal relationship required for robust generalization (Xing et al., 2025).
@@
A natural solution for OOD tasks is Supervised Fine-Tuning (SFT) with task-specific data. However, SFT approaches based on imitation learning rely heavily on costly, high-quality human demonstrations. When only a limited number of demonstrations are available, their effectiveness drops sharply. In contrast, online Reinforcement Learning (RL) provides an appealing alternative by enabling autonomous data collection and policy refinement in deployment (Lu et al., 2025; Liu et al., 2025; Shu et al., 2025; Chen et al., 2025c; Tan et al., 2025a). However, while recent online RL-based post-training methods enable autonomous adaptation without requiring large amounts of expert demonstrations, they are not inherently designed to guarantee robustness against environmental perturbations. General RL fine-tuning focuses on maximizing task-specific reward signals in nominal environments, but does not explicitly regularize the sensitivity of the VLA model to environmental noise. Such optimized models are often over-adapted to the specific dynamics of the post-training environment and become brittle to even slight perturbations. This limitation highlights the need for principled approaches that not only optimize for task performance, but also explicitly constrain the sensitivity of the model’s decision process to perturbations. Such robustness-oriented post-training is crucial for ensuring reliable deployment of VLA models in real-world environments.
@@
Environmental perturbations can arise from multiple sources and affect different stages of the model-environment interaction (Gu et al., 2025). Currently, we focus on two fundamental types: observation perturbations and action perturbations. Observation perturbations reflect inaccuracies in perception, caused by sensor noise, latency, or imperfect state estimation, which lead to mismatches between the perceived and actual environment state. Action perturbations, on the other hand, arise from actuation errors, implementation inaccuracies, or hardware-level disturbances, which cause deviations between the intended and executed actions. These two perturbations, which are ubiquitous in real-world robotic systems, can be effectively modeled as random noise applied to the sequential decision-making process (Duan et al., 2016). Furthermore, if not properly addressed, these perturbations can amplify over time and significantly degrade model performance. Thus, a significant gap remains: ensuring the robustness of post-training to environmental perturbations, so that the resulting VLA model is stable and reliable in autonomous interactions.
@@ 2 Related Work
Recently, a growing body of research has explored RL as a mechanism to adapt VLA models to complex embodied tasks. The first research direction emphasizes offline RL training, where large-scale policy optimization can be performed without expensive online interactions. Previous work proposed the Q-Transformer (Chebotar et al., 2023), introducing autoregressive offline Q-learning with a Transformer. Subsequently, in the work by ReinboT (Zhang et al., 2025c), researchers proposed an end-to-end VLA model that incorporates the RL principle of maximizing cumulative return. Further work has proposed an offline RL post-training algorithm, ARFM (Zhang et al., 2025b), for VLA flow models to boost the performance of VLA models on downstream tasks. Moreover, some work has focused on online RL fine-tuning of VLA models. FLaRe (Hu et al., 2024) employs large-scale domain randomized RL fine-tuning, and PA-RL (Mark et al., 2024) unifies offline and online RL across policy categories. IRe-VLA (Guo et al., 2025), RIPT-VLA (Tan et al., 2025a), and VLA-RL (Lu et al., 2025) are specifically designed to stabilize or extend online RL training of VLA models. Furthermore, optimization advances such as TGRPO (Chen et al., 2025c) and RFTF (Shu et al., 2025) improve the efficiency of fine-tuning in trajectory or sparse reward settings. RLDG (Xu et al., 2024) refines task-specific RL into a general VLA, ConRFT (Chen et al., 2025b) combines offline Q-learning with consistency-based online fine-tuning, and ReWiND (Zhang et al., 2025d) introduces a language-guided reward model for fast task adaptation without the need for new demonstrations. Moreover, several works have focused on training value functions for visual-language manipulation tasks to align VLA models with downstream tasks, including Bellman-Guided Retrials (Du et al., 2024), V-GPS (Nakamoto et al., 2024), and Hume (Song et al., 2025) However, previous work has often neglected the robustness and reliability of robotic visual-language manipulation under environmental perturbations and distribution changes, which is precisely the focus of our work. More related works are in Appendix A.1.
@@
Robust RL generally studies how to ensure stability and generalization under distribution shifts, noisy observations, and adversarial perturbations. Classical approaches leverage robust MDP formulations (Iyengar, 2005; Wiesemann et al., 2013) and adversarial training (Pinto et al., 2017). Modern deep RL has introduced diverse techniques, including domain randomization for sim-to-real transfer (Tobin et al., 2017; Van et al., 2025) and methods for offline and curriculum learning (Dennis et al., 2020; Wu et al., 2024). Recently, the field has advanced towards theoretically-grounded distributionally robust optimization (Li et al., 2024; Cui et al., 2025; He et al., ) with provable finite-sample guarantees (Ghosh et al., 2025; Roch et al., 2025). However, the online setting where an agent learns via direct interaction is less studied. A key challenge, unaddressed by previous work that often assumes access to exploratory policies (Wang & Zou, 2021; Badrinath & Kalathil, 2021), is to handle exploration and robustness simultaneously. Crucially, this focus on test-time robustness to environmental shifts distinguishes distributionally robust RL from corruption-robust RL, which handles corrupted training data (Lykouris et al., 2021; Wei et al., 2022; Zhang et al., 2022; Ye et al., 2024; 2023). Different from these, our work focuses on the post-training of VLA models for general task settings, aiming to simultaneously address model adaptability and stability issues within a unified model framework.
@@ 3 Preliminaries
RL post-training of VLA model. We model the task as a Markov Decision Process (MDP) (Sutton et al., 1998). The VLA model $\pi_{\theta}$ maps a natural language instruction $g$ and a sequence of observations $o_{1:t}$ and previous actions $a_{1:t-1}$ to a probability distribution over the current action $a_{t}$. Each episode is initialized with the context $\mathbf{c}=\left(g,o_{1}\right)$. The state $s_{t}$ is represented as $\left[g,o_{1:t},a_{1:t-1}\right]$. At each time step $t$, $\pi_{\theta}$ samples an action from the model distribution: $a_{t}\sim\pi_{\theta}\left(\cdot\mid g,o_{1:t},a_{1:t-1}\right)$. The environment $\mathcal{E}$ transitions to the next state $s_{t+1}=\left[g,o_{1:t+1},a_{1:t}\right]$. The environment $\mathcal{E}$ also returns a binary reward $R=1$ if the task goal is achieved, and $R=0$ otherwise. The goal of $\pi_{\theta}$ is to maximize the expected task reward: $R_{\theta}(\mathbf{c})=\mathbb{E}_{\mathbf{a}\sim\pi_{\theta}(\cdot\mid\mathbf{c})}[R(\mathbf{c},\mathbf{a})].$ To optimize this objective, policy gradient (Sutton et al., 1999) is a simple and efficient method: $\nabla_{\theta}\mathcal{L}_{\theta}(\mathbf{c})=\mathbb{E}_{\mathbf{a}\sim\pi_{\theta}}\left[\nabla_{\theta}\log\pi_{\theta}(\mathbf{a}\mid\mathbf{c})A(\mathbf{c},\mathbf{a})\right],$ where $A(\mathbf{c},\mathbf{a})$ is the advantage function, which indicates how good the action $\mathbf{a}$ is compared to the baseline. However, the accurate estimation of $A(\mathbf{c},\mathbf{a})$ is tricky, especially in the post-training phase of VLA models. To alleviate this issue, recent work (Chen et al., 2025a) proposed a critic-free optimization framework. Specifically, for each sampling context $\mathbf{c}$, we can draw $K$ rollouts $\left\{\mathbf{a}_{k}\sim\pi_{\psi}(\cdot\mid\mathbf{c})\right\}_{k=1}^{K}$ under a sampling model $\pi_{\psi}$. Each rollout receives a reward $R_{k}=R\left(\mathbf{c},\mathbf{a}_{k}\right)$. The leave-one-out baseline for rollout $k$ is calculated by averaging the other rewards (Kool et al., 2018): $A_{k}=R_{k}-\frac{1}{K-1}\sum_{j\neq k}R_{j}.$ This allows us to bypass the cumbersome task of learning a value function and efficiently compute a stable advantage signal. Furthermore, to utilize the collected rollouts $\left\{\left(\mathbf{c}_{k},\mathbf{a}_{k},A_{k}\right)\right\}$ to update $\pi_{\theta}$, we compute the importance ratio: $r_{k}=\pi_{\theta}\left(\mathbf{a}_{k}\mid\mathbf{c}_{k}\right)/\pi_{\psi}\left(\mathbf{a}_{k}\mid\mathbf{c}_{k}\right)$, where $\pi_{\theta}$ is the current update model and $\pi_{\psi}$ is a sampling model. We then optimize $\pi_{\theta}$ utilizing the Proximal Policy Optimization (PPO) objective (Schulman et al., 2017):
$\mathcal{L}_{\mathrm{PPO}}=-\min\left(r_{i}A_{i},\operatorname{clip}\left(r_{i},1-\epsilon,1+\epsilon\right)A_{i}\right),$ (1)
@@
Environmental Perturbations During RL Post-Training. We formalize the robust post-training problem of VLA model $\pi_{\theta}$ under environmental perturbations. At each time step $t$ in MDP, the $\pi_{\theta}$ receives a perturbed observation $\tilde{s}_{t}$ of the true state $s_{t}$, with deviation bounded by $\|\tilde{s}_{t}-s_{t}\|\leq\epsilon_{s}.$ The action selected by the $\pi_{\theta}$ may be corrupted by execution noise, modeled as Gaussian perturbations: $a_{t}=\pi_{t}(\tilde{s}_{t})+\xi_{t}$, where $\xi_{t}\sim\mathcal{N}(0,\sigma^{2}I_{d})$. The environment evolves according to a Lipschitz continuous transition function $f$, ensuring that small deviations in states or actions yield bounded deviations in subsequent states. Finally, we denote by $\epsilon_{\text{offline}}$ the discrepancy between the initial VLA model $\pi_{init}$, obtained via imitation learning, and the expert model $\pi^{*}$. This discrepancy provides the starting error baseline for the online post-training process.
@@ 4.1 Robustness analysis to environmental perturbations
In this section, we theoretically quantify how different types of perturbations affect the robustness of the VLA model. Theorem 1 discusses the return error amplification bound under observation perturbations and provides its specific relationship to the model’s Jacobian sensitivity. Theorem 2 focuses on return drift control under action noise, exploring the relationship between the return gap and the cumulative model drift. Furthermore, Theorem 3 considers the robust stability guarantee in combined settings involving both observation and action perturbations.
@@
$J(\pi^{*})-J(\pi)\leq\mathcal{O}(HL_{r}L_{f}^{H})\cdot\left(\epsilon_{\text{offline}}+\lambda\epsilon_{s}\right).$
Theorem 1 indicates that the robustness gap under observation perturbations is governed by the product $\lambda\epsilon_{s}$, where $\lambda$ measures the VLA model’s local sensitivity and $\epsilon_{s}$ quantifies the observation noise level. This means that even when the environment dynamics are stable ($L_{f}<1$), a large $\lambda$ can amplify small perturbations, leading to significant performance degradation. Conversely, enforcing Jacobian regularization to shrink $\lambda$ directly reduces the error propagation term $\lambda\epsilon_{s}$, thereby bounding $J(\pi^{*})-J(\pi)$ more tightly.
@@
Assume actions are perturbed as $a_{t}=\pi_{t}(s_{t})+\xi_{t}$ with $\xi_{t}\sim\mathcal{N}(0,\sigma^{2}I_{d})$, and VLA models satisfy $\|\pi_{t}-\pi_{t-1}\|_{\infty}\leq\delta_{t}$. Then:
$J(\pi^{*})-J(\pi)\leq\mathcal{O}(HL_{r}L_{f}^{H})\cdot\left(\epsilon_{\text{offline}}+\sum_{t=1}^{H}\delta_{t}+\sigma\sqrt{d}\right).$
Theorem 2 reveals that the return gap under action perturbations has two additive drivers: 1) the cumulative VLA model drift $\sum_{t=1}^{H}\delta_{t}$, which reflects the nonstationarity of online updates, and 2) the stochastic execution noise term $\sigma\sqrt{d}$, which is irreducible and scales with action dimension $d$. The presence of $\sum_{t=1}^{H}\delta_{t}$ shows that if VLA model updates are too aggressive (large $\delta_{t}$), then even in the absence of large execution noise ($\sigma\approx 0$), the compounding drift will destabilize rollouts and inflate $J(\pi^{*})-J(\pi)$. This means smoothness regularization—enforcing $\delta_{t}$ to be small—is essential to decouple online learning progress from robustness degradation, ensuring that long-horizon performance does not deteriorate due to accumulated VLA model shifts.
@@
$J(\pi^{*})-J(\pi)\leq\mathcal{O}(H^{2}L_{r}L_{f}^{H})\cdot\left(\epsilon_{\text{offline}}+\sum_{t=1}^{H}\delta_{t}+\lambda\epsilon_{s}+\sigma\sqrt{d}\right).$
Theorem 3 shows that when both observation and action perturbations are present, the return gap depends jointly on $\lambda\epsilon_{s}$ and $\sigma\sqrt{d}$, but with an amplified scaling factor of $\mathcal{O}(H^{2}L_{f}^{H})$. The quadratic dependence on $H$ arises because state deviations caused by noisy inputs ($\lambda\epsilon_{s}$) not only affect immediate rewards but also feed into subsequent actions, which themselves are perturbed by $\sigma\sqrt{d}$ and policy drift $\sum_{t=1}^{H}\delta_{t}$. This nested feedback loop means that robustness cannot be achieved by controlling only one source of error. Instead, both Jacobian regularization (reducing $\lambda$) and smoothness regularization (limiting $\delta_{t}$) must be applied simultaneously to keep the compounded gap bounded. This finding emphasizes that robust RL fine-tuning requires a dual regularization strategy, as ignoring either perturbation channel will inevitably result in superlinear error accumulation. More robustness analysis is in the Appendix A.2.

### v2

@@ 1 Introduction
Vision-Language-Action (VLA) models have emerged as a powerful paradigm for general-purpose robotic manipulation, with representative examples including RT-1 (brohan2022rt), RT-2 (zitkovich2023rt), OpenVLA (kim2024openvla), and $\pi_{0}$ (black2024pi_0). These VLA models demonstrate the feasibility of learning universal robotic policies from large-scale multimodal datasets and have achieved remarkable performance across a variety of manipulation tasks. Despite these advances, recent studies (xing2025shortcut; shi2025diversity) reveal that the generalization ability of pre-trained VLA models remains limited when deployed in Out-of-Distribution (OOD) scenarios. The OOD scenarios in this work primarily refer to variations in the distribution of perception and execution levels. These variations are caused by unseen visual conditions (lighting, occlusion, camera changes, etc.) and execution disturbances (action noise).
One key reason is that VLA models may rely on spurious correlations between irrelevant observation features and actions, rather than capturing the true causal relationship required for robust generalization (xing2025shortcut).
@@
A natural solution for OOD tasks is Supervised Fine-Tuning (SFT) with task-specific data. However, SFT approaches based on imitation learning rely heavily on costly, high-quality human demonstrations. When only a limited number of demonstrations are available, their effectiveness drops sharply. In contrast, online Reinforcement Learning (RL) provides an appealing alternative by enabling autonomous data collection and policy refinement in deployment (lu2025vla; liu2025can; shu2025rftf; chen2025tgrpo; tan2025interactive). However, while recent online RL-based post-training methods enable autonomous adaptation without requiring large amounts of expert demonstrations, they are not inherently designed to guarantee robustness against environmental perturbations. General RL fine-tuning focuses on maximizing task-specific reward signals in nominal environments, but does not explicitly regularize the sensitivity of the VLA model to environmental noise. Such optimized models are often over-adapted to the specific dynamics of the post-training environment and become brittle to even slight perturbations. This limitation highlights the need for principled approaches that not only optimize for task performance, but also explicitly constrain the sensitivity of the model’s decision process to perturbations. Such robustness-oriented post-training is crucial for ensuring reliable deployment of VLA models in real-world environments.
@@
Environmental perturbations can arise from multiple sources and affect different stages of the model-environment interaction (gu2025robust). Currently, we focus on two fundamental types: observation perturbations and action perturbations. Observation perturbations reflect inaccuracies in perception, caused by sensor noise, latency, or imperfect state estimation, which lead to mismatches between the perceived and actual environment state. Action perturbations, on the other hand, arise from actuation errors, implementation inaccuracies, or hardware-level disturbances, which cause deviations between the intended and executed actions. These two perturbations, which are ubiquitous in real-world robotic systems, can be effectively modeled as random noise applied to the sequential decision-making process (duan2016benchmarking). Furthermore, if not properly addressed, these perturbations can amplify over time and significantly degrade model performance. Thus, a significant gap remains: ensuring the robustness of post-training to environmental perturbations, so that the resulting VLA model is stable and reliable in autonomous interactions.
@@ 2 Related Work
Recently, a growing body of research has explored RL as a mechanism to adapt VLA models to complex embodied tasks. The first research direction emphasizes offline RL training, where large-scale policy optimization can be performed without expensive online interactions. Previous work proposed the Q-Transformer (chebotar2023q), introducing autoregressive offline Q-learning with a Transformer. Subsequently, in the work by ReinboT (zhang2025reinbot), researchers proposed an end-to-end VLA model that incorporates the RL principle of maximizing cumulative return. Further work has proposed an offline RL post-training algorithm, ARFM (zhang2025balancingsignalvarianceadaptive), for VLA flow models to boost the performance of VLA models on downstream tasks. Moreover, some work has focused on online RL fine-tuning of VLA models. FLaRe (hu2024flare) employs large-scale domain randomized RL fine-tuning, and PA-RL (mark2024policy) unifies offline and online RL across policy categories. IRe-VLA (guo2025improving), RIPT-VLA (tan2025interactive), and VLA-RL (lu2025vla) are specifically designed to stabilize or extend online RL training of VLA models. Furthermore, optimization advances such as TGRPO (chen2025tgrpo) and RFTF (shu2025rftf) improve the efficiency of fine-tuning in trajectory or sparse reward settings. RLDG (xu2024rldg) refines task-specific RL into a general VLA, ConRFT (chen2025conrft) combines offline Q-learning with consistency-based online fine-tuning, and ReWiND (zhang2025rewind) introduces a language-guided reward model for fast task adaptation without the need for new demonstrations. Moreover, several works have focused on training value functions for visual-language manipulation tasks to align VLA models with downstream tasks, including Bellman-Guided Retrials (du2024err), V-GPS (nakamoto2024steering), and Hume (song2025hume) However, previous work has often neglected the robustness and reliability of robotic visual-language manipulation under environmental perturbations and distribution changes, which is precisely the focus of our work. More related works are in Appendix A.1.
@@
Robust RL generally studies how to ensure stability and generalization under distribution shifts, noisy observations, and adversarial perturbations. Classical approaches leverage robust MDP formulations (iyengar2005robust; wiesemann2013robust) and adversarial training (pinto2017robust). Modern deep RL has introduced diverse techniques, including domain randomization for sim-to-real transfer (tobin2017domain; van2025hybrid) and methods for offline and curriculum learning (dennis2020emergent; wu2024acl). Recently, the field has advanced towards theoretically-grounded distributionally robust optimization (li2024safe; cui2025dr; hesample) with provable finite-sample guarantees (ghosh2025provably; roch2025finite). However, the online setting where an agent learns via direct interaction is less studied. A key challenge, unaddressed by previous work that often assumes access to exploratory policies (wang2021online; badrinath2021robust), is to handle exploration and robustness simultaneously. Crucially, this focus on test-time robustness to environmental shifts distinguishes distributionally robust RL from corruption-robust RL, which handles corrupted training data (lykouris2021corruption; wei2022model; zhang2022corruption; ye2024towards; ye2023corruption). Different from these, our work focuses on the post-training of VLA models for general task settings, aiming to simultaneously address model adaptability and stability issues within a unified model framework.
@@ 3 Preliminaries
RL post-training of VLA model. We model the task as a Markov Decision Process (MDP) (sutton1998reinforcement). The VLA model $\pi_{\theta}$ maps a natural language instruction $g$ and a sequence of observations $o_{1:t}$ and previous actions $a_{1:t-1}$ to a probability distribution over the current action $a_{t}$. Each episode is initialized with the context $\mathbf{c}=\left(g,o_{1}\right)$. The state $s_{t}$ is represented as $\left[g,o_{1:t},a_{1:t-1}\right]$. At each time step $t$, $\pi_{\theta}$ samples an action from the model distribution: $a_{t}\sim\pi_{\theta}\left(\cdot\mid g,o_{1:t},a_{1:t-1}\right)$. The environment $\mathcal{E}$ transitions to the next state $s_{t+1}=\left[g,o_{1:t+1},a_{1:t}\right]$. The environment $\mathcal{E}$ also returns a binary reward $r_{t}=1$ if the task goal is achieved, and $r_{t}=0$ otherwise. The goal of $\pi_{\theta}$ is to maximize the undiscounted task reward: $J(\pi_{\theta})=\mathbb{E}_{\tau\sim\pi_{\theta}}[\sum_{t=1}^{H}r_{t}],$ where $H$ is the rollout horizon. To optimize this objective, policy gradient (sutton1999policy) is a simple and efficient method: $\nabla_{\theta}J(\pi_{\theta})=\mathbb{E}_{\mathbf{a}\sim\pi_{\theta}}\left[\nabla_{\theta}\log\pi_{\theta}(\mathbf{a}\mid\mathbf{c})A(\mathbf{c},\mathbf{a})\right],$ where $A(\mathbf{c},\mathbf{a})$ is the advantage function, which indicates how good the action $\mathbf{a}$ is compared to the baseline. However, the accurate estimation of $A(\mathbf{c},\mathbf{a})$ is tricky, especially in the post-training phase of VLA models. To alleviate this issue, recent work (chen2025reinforcement) proposed a critic-free optimization framework. Specifically, for each sampling context $\mathbf{c}$, we can draw $K$ rollouts $\left\{\mathbf{a}_{k}\sim\pi_{\psi}(\cdot\mid\mathbf{c})\right\}_{k=1}^{K}$ under a sampling model $\pi_{\psi}$. Each rollout receives a reward $r_{k}=r\left(\mathbf{c},\mathbf{a}_{k}\right)$. The leave-one-out baseline for rollout $k$ is calculated by averaging the other rewards (kool2018attention): $A_{k}=r_{k}-\frac{1}{K-1}\sum_{j\neq k}r_{j}.$ This allows us to bypass the cumbersome task of learning a value function and efficiently compute a stable advantage signal. Furthermore, to utilize the collected rollouts $\left\{\left(\mathbf{c}_{k},\mathbf{a}_{k},A_{k}\right)\right\}$ to update $\pi_{\theta}$, we compute the importance ratio: $\eta_{k}=\pi_{\theta}\left(\mathbf{a}_{k}\mid\mathbf{c}_{k}\right)/\pi_{\psi}\left(\mathbf{a}_{k}\mid\mathbf{c}_{k}\right)$, where $\pi_{\theta}$ is the current update model and $\pi_{\psi}$ is a sampling model. We then optimize $\pi_{\theta}$ utilizing the Proximal Policy Optimization (PPO) objective (schulman2017proximal):
$\mathcal{L}_{\mathrm{PPO}}=-\min\left(\eta_{i}A_{i},\operatorname{clip}\left(\eta_{i},1-\epsilon,1+\epsilon\right)A_{i}\right),$ (1)
@@
Environmental Perturbations During RL Post-Training. We formalize the robust post-training problem of VLA model $\pi_{\theta}$ under environmental perturbations. At each time step $t$ in MDP, the $\pi_{\theta}$ receives a perturbed observation $\tilde{s}_{t}$ of the true state $s_{t}$, with deviation bounded by $\|\tilde{s}_{t}-s_{t}\|\leq\epsilon_{s}.$ The action selected by the $\pi_{\theta}$ may be corrupted by execution noise, modeled as Gaussian perturbations: $a_{t}=\pi_{t}(\tilde{s}_{t})+\xi_{t}$, where $\xi_{t}\sim\mathcal{N}(0,\sigma^{2}I_{d})$. The environment evolves according to a Lipschitz continuous transition function $f$, ensuring that small deviations in states or actions yield bounded deviations in subsequent states. The environmental dynamics $f(s,a)$ and the reward function $r(s,a)$ are Lipschitz continuous and are constants $L_{f}$ and $L_{r}$, respectively. Finally, we denote by $\epsilon_{\text{offline}}$ the discrepancy between the initial VLA model $\pi_{init}$, obtained via imitation learning, and the expert model $\pi^{*}$. This discrepancy provides the starting error baseline for the online post-training process.
@@ 4.1 Robustness analysis to environmental perturbations
In the perturbed online RL post-training of the VLA model, we observed two main sources of robustness degradation. First, these models are highly sensitive to small changes in visual input: slight shifts in lighting, occlusions, or camera noise can cause disproportionately large changes in the model’s latent representation and therefore in its actions. This motivates explicitly constraining how sharply the model reacts to observation changes. Second, when the model is updated online, its behavior can change abruptly from one iteration to the next. Even small gradient steps may lead to large jumps in the produced actions, which become unstable when combined with the stochasticity of realistic execution. To stabilize this process, we also need to control how quickly the model is allowed to evolve during post-training. Therefore, in this section, we theoretically quantify how different types of perturbations affect the robustness of the VLA model.
@@
$\mathbb{E}\big[J(\pi^{*})-J(\pi)\big]\leq\mathcal{O}(HL_{r}L_{f}^{H})\cdot\left(\epsilon_{\text{offline}}+\lambda\epsilon_{s}\right).$}
Here the factor $L_{f}^{H}$ corresponds to the worst-case geometric amplification under $L_{f}$-Lipschitz dynamics; when $L_{f}<1$ the bound becomes strictly smaller. Theorem 1 indicates that the robustness gap under observation perturbations is governed by the product $\lambda\epsilon_{s}$, where $\lambda$ measures the VLA model’s local sensitivity and $\epsilon_{s}$ quantifies the observation noise level. Conversely, enforcing Jacobian regularization to shrink $\lambda$ directly reduces the error propagation term $\lambda\epsilon_{s}$, thereby bounding $J(\pi^{*})-J(\pi)$ more tightly.
@@
Assume actions are perturbed as $a_{t}=\pi_{t}(s_{t})+\xi_{t}$ with $\xi_{t}\sim\mathcal{N}(0,\sigma^{2}I_{d})$, and VLA models satisfy $\|\pi_{i}-\pi_{i-1}\|_{\infty}\leq\delta_{i}$}. Then:
$\mathbb{E}\big[J(\pi^{*})-J(\pi)\big]\leq\mathcal{O}(HL_{r}L_{f}^{H})\cdot\Big(\epsilon_{\mathrm{offline}}+{\sum_{i=1}^{N}\delta_{i}}+\sigma\sqrt{d}\Big).$}
Here, $N$ represents the number of times the model parameters are updated. Theorem 2 reveals that the return gap under action perturbations has two additive drivers: 1) the cumulative VLA model drift $\sum_{i=1}^{N}\delta_{i}$}, which reflects the nonstationarity of online updates, and 2) the stochastic execution noise term $\sigma\sqrt{d}$, which is irreducible and scales with action dimension $d$. The presence of $\sum_{i=1}^{N}\delta_{i}$} shows that if VLA model updates are too aggressive (large $\delta_{i}$}), then even in the absence of large execution noise ($\sigma\approx 0$), the compounding drift will destabilize rollouts and inflate $J(\pi^{*})-J(\pi)$. This means smoothness regularization -- enforcing $\delta_{i}$} to be small -- is essential to decouple online learning progress from robustness degradation, ensuring that long-horizon performance does not deteriorate due to accumulated VLA model shifts.
@@
$\mathbb{E}\big[J(\pi^{*})-J(\pi)\big]\leq\mathcal{O}(HL_{r}L_{f}^{H})\cdot\left(\epsilon_{\text{offline}}+\sum_{i=1}^{N}\delta_{i}+\lambda\epsilon_{s}+\sigma\sqrt{d}\right).$}
Theorem 3 shows that when both observation and action perturbations are present, the return gap depends jointly on the observation–sensitivity term $\lambda\epsilon_{s}$, and the accumulated update drift $\sum_{i=1}^{N}\delta_{i}$. Unlike the single-perturbation cases, here the deviation caused by noisy observations does not remain localized: it propagates through the dynamics, influences later states, and interacts with execution noise and model drift across the entire rollout. This nested feedback structure implies that robustness cannot be achieved by controlling only one source of error. Instead, Jacobian regularization (which reduces the instantaneous sensitivity $\lambda$) and smoothness regularization (which limits update-induced drift $\delta_{i}$) must be applied jointly to prevent compounding error growth over long horizons. This result highlights that robust VLA post-training fundamentally requires a dual regularization strategy; ignoring either perturbation channel will allow errors to accumulate through the dynamics and ultimately destabilize the VLA model. More robustness analysis is in the Appendix A.2.